Given a sample size of 25, what would be the t-multiplier for a 95%, two-sided, confidence interval on mu?
Answer #1

sample size, n = 25
degrees of freedom, df = n - 1 = 24

Given CI level is 95%, hence α = 1 - 0.95 = 0.05
α/2 = 0.05/2 = 0.025, tc = t(α/2, df) = 2.064